OpenStudy (anonymous):

Tua has created a new game called Making Green. To play the game, a player spins twice. On one spinner, it has blue,red,yellow, and orange.(equally spilt into four) The other spinner has red,blue, orange, and yellow.(red,blue,orange equally spilt into three. Yellow is half the circle.) If the player gets blue in one section and yellow in the other, the player wins, because blue and yellow together make green. The player can choose either spinner for either of their two spins. Is there a different choice of selecting the spinners to make the probability of making green greater than that in part

OpenStudy (dumbcow):

3 possible choices: AA AB BB For AA, he uses spinner A for both spins P(green) = (1/4*1/4) + (1/4*1/4) = 2/16 = 1/8 For AB, 1 spin uses spinner A, the other spin with spinner B P(green) = (1/4*1/2) + (1/4*1/6) = 4/24 = 1/6 For BB, spinner B used in both spins P(green) = (1/2*1/6) + (1/6*1/2) = 2/12 = 1/6 Best chance at making green is using both spinners or only spinner B